# A microscopic theory of Anderson localization of electrons in random lattices

### v1

@@ III.4 High-dimensional approximation
It isn’t easy to solve the parquet equations (29) or (30) without approximations. It is more important to know whether these equations can lead to a critical point and a transition to a localized state than to solve them numerically exactly. We hence resort to an approximation justified in the asymptotic limit to infinite spatial dimension with a control parameter $d^{-1}$. A local approximation for the irreducible vertex $\bar{\Lambda}$ will guarantee the accuracy of the approximation to leading order in $d^{-1}$.
@@
In the next step, we close the equation for the local vertex $\bar{\Lambda}_{0}(E;\omega)$. We use identity $\bar{\Lambda}^{ee}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})=\bar{\Lambda}^{eh}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$. The solution of Eq. (31) leads to a non-local electron-hole irreducible vertex. The self-consistent equation for the local vertex $\bar{\Lambda}_{0}(E;\omega)$ is obtained as a projection of Eq. (31) to a spatially local form
@@
in this fully self-consistent solution. We denoted $\left\langle f(\mathbf{q})\right\rangle_{\mathbf{q}}\equiv N^{-1}\sum_{\mathbf{q}}f(\mathbf{q})$, summation over the transfer wave vectors. The self-consistent solution guarantees the existence of the diffusion pole at $\omega=0$, since $\Lambda_{0}(E;0)\chi(E;\mathbf{0})=1$ and similarly $\bar{\Lambda}_{0}(E;0)\bar{\chi}(E;\mathbf{0})=1$, as a direct consequence of Eqs. (33) and (34).
@@
This dynamical local approximation cannot be directly continued below the upper critical dimension $d_{u}=4$. The wave vector integrals on both sides of Eq. (36) diverge for $d\leq 4$. We can, however, keep the local irreducible vertex and the electron-hole bubble static in the denominator of the integrand on the right-hand side of Eq. (32). In that case, we can continue the equation
@@ III.5 Mean-field critical behavior in dimensions $d>2$
The local approximation of the electron-hole irreducible vertex is justified in high spatial dimensions, suppressing wave vector convolutions. The small parameter in the limit $d\to\infty$ in Eqs (37) is the bubble $\bar{\chi}(\mathbf{q})$. We expand the integrand in powers of this function to simplify the self-consistent equation for the local vertex $\Lambda_{0}(E;\omega)$. We obtain in the leading order a cubic equation
@@
The qualitative, mean-field-like behavior will be guaranteed if we replace vertex $\Lambda_{0}(E;0)$ in the denominator of the second term on the right-hand side of Eq. (37) by the CPA irreducible vertex $\lambda(E;0)$ and the CPA bubble ${\chi}_{0}(E;0,\mathbf{q})$ so that to maintain the diffusion pole. In this simplification, we do not alter the qualitative behavior of vertex $\Lambda_{0}(E;\omega)$ in $d>4$. The irreducible vertex is then determined from a quadratic equation
@@
This solution leads to a bifurcation point at $4a(E;0)\lambda_{c}=1$ at which it splits into two complex conjugate roots. It is also immediately apparent that $a(E;0)\to 0$ and $\lambda_{c}\to\infty$ when $d\to\infty$. The upper critical dimension for vanishing of this bifurcation point is then $d=\infty$, where the CPA becomes exact. For the disorder strength $\lambda(E;0)>\lambda_{c}$ the irreducible vertex becomes complex at zero transfer energy $\omega\to 0$ [1]
@@
The irreducible vertex $\Lambda_{0}(E;\omega)=\Lambda_{R}(E;\omega)+i\Lambda_{I}(E;\omega)$ becomes discontinuous at $\omega=0$ with a jump in the imaginary part $\Lambda_{I}(E;\omega)$. The order parameter in the new phase is $|\Lambda_{I}(E;0)|$ and its critical exponent in the local approximation is
@@ IV Anderson localization
We have dealt so far with the renormalized perturbation theory for the one-particle self-energy and the two-particle irreducible vertices within and beyond the CPA. We found a critical point at which the frequency derivative of the electron-hole irreducible vertex diverges and a new phase emerges. We have not yet proven that this new state is indeed Anderson localization. We must relate the microscopically derived new phase to the macroscopically observed properties that characterize Anderson localization. The first step in this direction is to define the conserving irreducible vertex $L_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$, guaranteeing the macroscopic conservation laws and the proper low-energy asymptotics of the diffusion pole in the electron-hole correlation function.
@@ IV.1 Dynamical vertex, conserving vertex, and the Ward identity
A common feature of the microscopic theory of disordered and interacting electrons is that microscopic diagrammatic perturbation theory cannot guarantee the full dynamical extent of conservation laws. The dynamical Ward identity, Eq. (14), is in conflict with causality, demanding $\Lambda_{\mathbf{k}\mathbf{k}^{\prime}}(E;0,\mathbf{q})>0$ in the metallic phase [50]. Only the static form of this Ward identity used in Eq. (25) can be guaranteed by the causal perturbation theory. We then have to introduce two two-particle vertices. The dynamical electron-hole vertex from the microscopic perturbation theory is determined from the Bethe-Salpeter equation with the irreducible vertex from the perturbation theory $\Lambda^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$ with the conserving vector $\mathbf{q}$ from the electron-hole scattering channel
@@
The conserving electron-hole vertex $\mathcal{K}^{RA}_{{\bf k}{\bf k}^{\prime}}(E;\omega,{\bf q})$ is obtained from the Bethe-Salpeter equation (7) with the conserving irreducible vertex $L^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$. It is derived from the dynamical vertex $\Lambda^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$ as follows [51]
@@
where the angular brackets denote the normalized integration over the fermionic momenta $\mathbf{k}$ of the enclosed functions. The difference function $\Delta G_{\mathbf{k}}(E;\omega,\mathbf{q})$ was deined in Eq. (14). The correcting function is
$R_{\mathbf{k}}(E;\omega,\mathbf{q})=\frac{1}{N}\sum_{\mathbf{k}^{\prime}}\\ \left\{\left[\Lambda^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})-\ \Lambda^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E_{+};0_{+},\mathbf{0})\right]G^{R}_{\mathbf{k}^{\prime}}(E_{+})\right.\\ \left.-\ \left[\Lambda^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})-\ \Lambda^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E_{-};0_{-},\mathbf{0})\right]G^{A}_{\mathbf{k}^{\prime}}(E_{-})\right\}\,.$ (47)
In the metallic phase, $R_{\mathbf{k}}(E;0,\mathbf{0})=0$. When extending the conserving vertex into the new state beyond the bifurcation point, we keep this condition, and the two irreducible vertices are identical in the homogeneous and static limit,
@@
This condition is important, in particular in the localized phase where $\Lambda^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;0,\mathbf{0})$ acquires an imaginary component and $\Lambda^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{0})$ is no longer continuous at $\omega=0$. For the same reason, the Ward identity when continued beyond the critical point into the localized phase must be modified to
@@ IV.2 Local approximation
We now calculate the conserving vertex $\mathcal{K}^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$ for the local approximation of the dynamical irredicible vertex, ${\Lambda}_{0}(E;\omega)$ from Eq. (42). A slightly different non-local approximation with $1/d$ corrections was used in Ref. [1] with a qualitatively similar result. The advantage of the local approximation for the irreducible electron-hole vertex is the locality of the corresponding self-energy. We have
@@
$R(E;\omega)=\left[\Lambda_{0}(E;\omega)-\Lambda_{0}(E_{+};0_{+})\right]\left\langle G^{R}(E_{+})\right\rangle\\ -\ \left[\Lambda_{0}(E;\omega)-\Lambda_{0}(E_{-};0_{-})\right]\left\langle G^{A}(E_{-})\right\rangle\,$ (51a)
@@
$\sigma(E;\omega,\mathbf{q})=-\frac{ie^{2}\omega}{4\mathcal{D}(E;\omega,\mathbf{q})}\left.\nabla_{\mathbf{q}}^{2}\mathcal{D}(E;0,\mathbf{q})\right|_{q=0}\,.$ (58)
@@ IV.2.1 Metallic phase
$\dot{\Lambda}_{I}(E;0)\\ =-\left\langle\left\langle G^{R}_{\mathbf{k}_{+}}(E)G^{A}_{\mathbf{k}_{-}}(E)\Im\Delta G_{\mathbf{k}}(E;0,\mathbf{q})\right\rangle\bar{\chi}(E;0,\mathbf{q})\right\rangle_{q}\\ \times\frac{1-2a(E;0)\lambda(E;0)-\sqrt{1-4a(E;0)\lambda(E;0)}}{2a(E;0)^{2}\sqrt{1-4a(E;0)\lambda(E;0)}}\,$, (66c)
@@
$\Re\sigma(E;\omega)$ $=\frac{e^{2}|\Im\Sigma^{R}(E)|}{1+A(E)^{2}\omega^{2}}\left[\left\langle\Im G^{R}(E)^{2}\nabla\epsilon^{2}\right\rangle-A(E)\left(A(E)-\frac{2\dot{\Lambda}_{I}(E;0)\left\langle\Im G^{R}(E)\right\rangle\left\langle\Im G^{R}(E)^{3}\nabla\epsilon^{2}\right\rangle}{\Im\Sigma^{R}(E)^{2}\left\langle\Im G^{R}(E)^{2}\right\rangle}\right)\omega^{2}\right]\,$, (67a)
@@
$\Im\sigma(E;\omega)$ $=-\frac{e^{2}|\Im\Sigma^{R}(E)|}{1+A(E)^{2}\omega^{2}}\left[A(E)\left\langle\Im G^{R}(E)^{2}\nabla\epsilon^{2}\right\rangle+\frac{2\dot{\Lambda}_{I}(E;0)\left\langle\Im G^{R}(E)\right\rangle\left\langle\Im G^{R}(E)^{3}\nabla\epsilon^{2}\right\rangle}{\Im\Sigma^{R}(E)^{2}\left\langle\Im G^{R}(E)^{2}\right\rangle}\right]\omega$ (67b)
@@
It is clear from the form of the dynamical conductivity that the transition to Anderson localization cannot be discerned from the static conductivity that is standardly calculated from the Kubo formula. As discussed in Ref. [1], Anderson localization is a bound state of the propagating particle and the hole left at the origin. The static quantities are spatially averaged and do not remember the diffusing particle’s initial position. Due to conservation laws, the diffusion pole of the electron-hole correlation function must be expanded to second order in frequency to observe the critical behavior of the metallic state at the Anderson localization transition.
IV.2.2 Localized phase

### v4

@@ III.4 High-dimensional approximation
It isn’t easy to solve the parquet equations (29) or (30) without approximations. It is more important to know whether these equations can lead to a critical point and a transition to a localized state than to solve them exactly. We hence resort to an approximation justified in the asymptotic limit of infinite spatial dimension, with a control parameter $d^{-1}$. A local approximation for the irreducible vertex $\bar{\Lambda}$ will guarantee the accuracy of the approximation to leading order in $d^{-1}$.
@@
In the next step, we close the equation for the local vertex $\bar{\Lambda}_{0}(E;\omega)$. We use the identity $\bar{\Lambda}^{ee}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})=\bar{\Lambda}^{eh}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$. The solution of Eq. (31) yields a non-local electron-hole irreducible vertex. The self-consistent equation for the local vertex $\bar{\Lambda}_{0}(E;\omega)$ is obtained by projecting Eq. (31) onto a spatially local form.
@@
in this fully self-consistent solution, we denote $\left\langle f(\mathbf{q})\right\rangle_{\mathbf{q}}\equiv N^{-1}\sum_{\mathbf{q}}f(\mathbf{q})$, a summation over the transfer wave vectors. The self-consistent solution guarantees the existence of the diffusion pole at $\omega=0$, since $\Lambda_{0}(E;0)\chi(E;\mathbf{0})=1$ and similarly $\bar{\Lambda}_{0}(E;0)\bar{\chi}(E;\mathbf{0})=1$, as a direct consequence of Eqs. (33) and (34).
@@
This dynamical local approximation cannot be directly continued below the upper critical dimension $d_{u}=4$. The wave-vector integrals on both sides of Eq. (36) diverge for $d\leq 4$. However, we can keep the local irreducible vertex and the electron-hole bubble static in the denominator of the integrand on the right-hand side of Eq. (32). In that case, we can continue the equation
@@ III.5 Mean-field critical behavior in dimensions $d>2$
The local approximation to the electron-hole irreducible vertex is justified in high spatial dimensions, suppressing wave-vector convolutions. The small parameter in the $d\to\infty$ limit of Eqs. (37) is the bubble $\bar{\chi}(\mathbf{q})$. We expand the integrand in powers of this function to simplify the self-consistent equation for the local vertex $\Lambda_{0}(E;\omega)$. In leading order, we obtain a cubic equation
@@
The qualitative, mean-field-like behavior will be guaranteed if we replace the vertex $\Lambda_{0}(E;0)$ in the denominator of the second term on the right-hand side of Eq. (37) with the CPA irreducible vertex $\lambda(E;0)$ and the CPA bubble ${\chi}_{0}(E;0,\mathbf{q})$ to preserve the diffusion pole. In this simplification, we do not alter the qualitative behavior of the vertex $\Lambda_{0}(E;\omega)$ in $d>4$. The irreducible vertex is then determined from a quadratic equation.
@@
This solution leads to a bifurcation point at $4a(E;0)\lambda_{c}=1$, where it splits into two complex conjugate roots. It is also immediately apparent that $a(E;0)\to 0$ and $\lambda_{c}\to\infty$ as $d\to\infty$. The upper critical dimension for the vanishing of this bifurcation point is then $d=\infty$, where the CPA becomes exact. For the disorder strength $\lambda(E;0)>\lambda_{c}$ the irreducible vertex becomes complex at zero transfer energy $\omega\to 0$ [1]
@@
The irreducible vertex $\Lambda_{0}(E;\omega)=\Lambda_{R}(E;\omega)+i\Lambda_{I}(E;\omega)$ becomes discontinuous at $\omega=0$ with a jump in the imaginary part $\Lambda_{I}(E;\omega)$. The order parameter in the new phase is $|\Lambda_{I}(E;0)|$, and its critical exponent in the local approximation is
@@ IV Anderson localization
We have so far developed renormalized perturbation theory for the one-particle self-energy and the two-particle irreducible vertices within and beyond the CPA. We found a critical point at which the frequency derivative of the electron-hole irreducible vertex diverges and a new phase emerges. We have not yet proven that this new state is indeed Anderson localization. We must relate the microscopically derived new phase to the macroscopically observed properties that characterize Anderson localization. The first step in this direction is to define the conserving irreducible vertex $L_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$, which guarantees the macroscopic conservation laws and the proper low-energy asymptotics of the diffusion pole in the electron-hole correlation function.
@@ IV.1 Dynamical vertex, conserving vertex, and the Ward identity
A common feature of the microscopic theory of disordered and interacting electrons is that diagrammatic perturbation theory cannot guarantee the full dynamical extent of conservation laws. The dynamical Ward identity, Eq. (14), conflicts with causality, demanding $\Lambda_{\mathbf{k}\mathbf{k}^{\prime}}(E;0,\mathbf{q})>0$ in the metallic phase [50]. Only the static form of this Ward identity, used in Eq. (25), can be guaranteed by causal perturbation theory. We then have to introduce two two-particle vertices. The dynamical electron-hole vertex from microscopic perturbation theory is determined by the Bethe-Salpeter equation, with the irreducible vertex from perturbation theory $\Lambda^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$ and the conserving vector $\mathbf{q}$ from the electron-hole scattering channel.
@@
The conserving electron-hole vertex $\mathcal{K}^{RA}_{{\bf k}{\bf k}^{\prime}}(E;\omega,{\bf q})$ is obtained from the Bethe-Salpeter equation (7) using the conserving irreducible vertex $L^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$. It is derived from the dynamical vertex $\Lambda^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$ as follows [51]
@@
where the angular brackets denote the normalized integration over the fermionic momenta $\mathbf{k}$ of the enclosed functions. The difference function $\Delta G_{\mathbf{k}}(E;\omega,\mathbf{q})$ was defined in Eq. (14). The correcting function is
$R_{\mathbf{k}}(E;\omega,\mathbf{q})=\frac{1}{N}\sum_{\mathbf{k}^{\prime}}\left\{\left[\Lambda^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})\right.\right.\\ \left.\left.-\ \Lambda^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E_{+};\mathrm{sign}(\omega)0_{+},\mathbf{0})\right]G^{R}_{\mathbf{k}^{\prime}}(E_{+})+\ \left[\Lambda^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})\right.\right.\\ \left.\left.-\ \Lambda^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E_{-};\mathrm{sign}(\omega)0_{-},\mathbf{0})\right]G^{A}_{\mathbf{k}^{\prime}}(E_{-})\right\}\,$, (47)
where we denoted $0_{\pm}=\pm\lim_{\epsilon\to 0}|\epsilon|$. In the metallic phase, $R_{\mathbf{k}}(E;0,\mathbf{0})=0$. When extending the conserving vertex into the new state beyond the bifurcation point, we keep this condition, and the two irreducible vertices are identical in the homogeneous and static limit,
@@
This condition is important, particularly in the localized phase, where $\Lambda^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;0,\mathbf{0})$ acquires an imaginary component and $\Lambda^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{0})$ is no longer continuous at $\omega=0$. For the same reason, the Ward identity, when continued beyond the critical point into the localized phase, must be modified to
@@ IV.2 Local approximation
We now calculate the conserving vertex $\mathcal{K}^{RA}_{\mathbf{k}\mathbf{k}^{\prime}}(E;\omega,\mathbf{q})$ for the local approximation to the dynamical irreducible vertex, ${\Lambda}_{0}(E;\omega)$, from Eq. (42). A slightly different non-local approximation with $1/d$ corrections was used in Ref. [1], with a qualitatively similar result. The advantage of the local approximation for the irreducible electron-hole vertex is that the corresponding self-energy is local. We have
@@
$R(E;\omega)\\ =\left[\Lambda_{0}(E;\omega)-\Lambda_{0}(E_{+};\mathrm{sign}(\omega)0_{+})\right]\left\langle G^{R}(E_{+})\right\rangle\\ -\ \left[\Lambda_{0}(E;\omega)-\Lambda_{0}(E_{-};\mathrm{sign}(\omega)0_{-})\right]\left\langle G^{A}(E_{-})\right\rangle\,$ (51a)
@@
$\sigma(E;\omega,\mathbf{q})=-\frac{ie^{2}\omega\ n(E)}{2\mathcal{D}(E;\omega,\mathbf{q})}\left.\nabla_{\mathbf{q}}^{2}\mathcal{D}(E;0,\mathbf{q})\right|_{q=0}\,.$ (58)
@@ IV.2.1 Metallic phase
$\dot{\Lambda}_{I}(E;0)=-\lambda(E;0)\\ \times\left\langle\left\langle G^{R}_{\mathbf{k}_{+}}(E)G^{A}_{\mathbf{k}_{-}}(E)\Im\Delta G_{\mathbf{k}}(E;0,\mathbf{q})\right\rangle\bar{\chi}(E;0,\mathbf{q})\right\rangle_{q}\\ \times\frac{1-2a(E;0)\lambda(E;0)-\sqrt{1-4a(E;0)\lambda(E;0)}}{2a(E;0)^{2}\sqrt{1-4a(E;0)\lambda(E;0)}}\,$, (66c)
@@
$\Re\sigma(E;\omega)$ $=\frac{e^{2}}{1+A(E)^{2}\omega^{2}}\left[\left\langle\Im G^{R}(E)^{2}\nabla\epsilon^{2}\right\rangle-A(E)\left(A(E)-\frac{2\dot{\Lambda}_{I}(E;0)\left\langle\Im G^{R}(E)\right\rangle\left\langle\Im G^{R}(E)^{3}\nabla\epsilon^{2}\right\rangle}{\Im\Sigma^{R}(E)^{2}\left\langle\Im G^{R}(E)^{2}\right\rangle}\right)\omega^{2}\right]\,$, (67a)
@@
$\Im\sigma(E;\omega)$ $=-\frac{e^{2}}{1+A(E)^{2}\omega^{2}}\left[A(E)\left\langle\Im G^{R}(E)^{2}\nabla\epsilon^{2}\right\rangle+\frac{2\dot{\Lambda}_{I}(E;0)\left\langle\Im G^{R}(E)\right\rangle\left\langle\Im G^{R}(E)^{3}\nabla\epsilon^{2}\right\rangle}{\Im\Sigma^{R}(E)^{2}\left\langle\Im G^{R}(E)^{2}\right\rangle}\right]\omega$ (67b)
@@
It is clear from the form of the dynamical conductivity that the transition to Anderson localization cannot be discerned from the static conductivity, which is standardly calculated from the Kubo formula. As discussed in Ref. [1], Anderson localization is a bound state of the propagating particle and the hole left at the origin. The static quantities are spatially averaged and do not remember the diffusing particle’s initial position. Due to conservation laws, the diffusion pole of the electron-hole correlation function must be expanded to second order in frequency to observe the critical behavior of the metallic state at the Anderson localization transition.
IV.2.2 Mobility edge and instability line of the metallic phase
The metallic phase is restricted to the region where the argument in the square root on the right-hand side of Eq. (42) is positive. In the static case, it is a real number that passes through zero at the mobility edge, as discussed in Sec. III.5. It is a bifurcation point below which a real value of the vertex function splits into two complex conjugate roots. In analogy with classical phase transitions, a symmetry-breaking field for the Anderson localization transition, which allows one to circumvent the critical point, was proposed in Ref. [11] to be the frequency difference $\omega$ in the electron-hole correlation function $\Phi^{RA}(E;\omega,\mathbf{q})$. However, contrary to standard statistical models of phase transitions, the frequency $\omega$ does not remove the critical point of the Anderson localization transition. The dynamic critical point at $\omega\neq 0$ does not separate a metal from an insulator, but indicates the genesis of an order parameter in the Anderson localized phase in the static limit, $\omega=0$. The critical point for $\omega\neq 0$ is determined by a change in the sign of the derivative $\partial\Im\Phi^{RA}(E;\omega,\mathbf{0})/\partial\omega$. This change occurs in the mean-field approximation when the real part of the argument of the square root in the solution for the irreducible vertex $\Lambda(E;\omega,\mathbf{0})$ goes through zero. The critical line in the $(E,\omega)$-plane from the approximation of Eq. (42) is determined by
$4\Re\left[\lambda(E_{c};\omega_{c})^{2}\left(\left\langle\chi(E_{c};\omega_{c},\mathbf{q})^{2}\right\rangle_{\mathbf{q}}\right.\right.\\ \left.\left.\qquad-\left\langle\chi(E_{c};\omega_{c},\mathbf{q})\right\rangle_{\mathbf{q}}^{2}\right)\right]=1\,.$ (68)
This critical line separates the perturbative metallic regime from a semimetallic phase, where the perturbative expansion breaks down and the static limit cannot be reached without introducing an order parameter. The critical line $(E_{c,\omega_{c}}$ resembles the de Almeida-Thouless line in the mean-field theory of spin glasses in a magnetic field [54]. This means that the expansion in frequency $\omega$ around the static limit cannot be used beyond the mobility edge $E_{c}$ for $\omega=0$, that is, below the instability line $(E_{c},\omega_{c})$, determined from Eq. (68).
We assess the behavior at the instability line for small frequencies below the mobility edge $(E_{c},0)$. Using the solution for the electron-hole irreducible vertex from Eq. (42) along the instability line, we obtain
$\Lambda_{c}$ $=\frac{2\lambda_{c}\left[1+\left(1+i\right)\lambda_{c}\sqrt{2\Im\Delta\chi_{2}}\right]}{1-2\lambda_{c}\sqrt{2\Im\Delta\chi_{2}}+4\lambda_{c}^{2}\Im\Delta\chi_{2}}$ (69)
where we neglected the imaginary part of the mean-field vertex $\lambda_{c}=\lambda(E_{c};\omega_{c})>0$ and denoted $\Delta\chi_{2}=\left\langle\chi(E_{c};\omega_{c},\mathbf{q})^{2}\right\rangle_{\mathbf{q}}-\left\langle\chi(E_{c};\omega_{c},\mathbf{q})\right\rangle_{\mathbf{q}}^{2}$.
We use Eq. (58) to determine the dynamic conductivity along the instability line for small frequencies
$\sigma(\omega)$ $=\sigma_{0}\frac{\omega\left[\left(1-i\alpha_{c}\right)\omega+2\pi n\Im\Lambda_{c}\right]}{\left(\omega+2\pi n\Im\Lambda_{c}\right)^{2}+\alpha_{c}^{2}\omega^{2}}\,$, (70)
where $n$ is the particle density, $\alpha_{c}=\Im\Lambda(E_{c};\omega_{c})/\Re\Lambda(E_{c};\omega_{c})\equiv\Im\Lambda_{c}/\Re\Lambda_{c}$, and $\sigma_{0}$ is the static conductivity at the mobility edge, calculated on the metallic side from Eq. (67). We see that $\Im\Lambda_{c}$ is the characteristic parameter of the instability line of the metallic phase and generates the real order parameter in the static limit.
The expression for the dynamic conductivity, Eq. (70), holds only for small frequencies near the mobility edge. It qualitatively captures the differences in the behavior of the dynamic conductivity near the mobility edge $E_{c}$. The mean-field solution in the metallic phase, $\Im\Lambda(E,0)=0$, yields the Drude expression for the dynamic conductivity, with $\sigma(0)=\sigma_{0}>0$ at the mobility edge. As we approach the mobility edge along the instability line, with $\Im\Lambda_{c}\propto|\omega|^{1/{2}}$, the static conductivity vanishes with a mean-field critical exponent $s=1/2$, that is, $\sigma(\omega)\propto\sqrt{|\omega|}$. Finally, the static conductivity below the mobility edge vanishes with a critical exponent $s=1$, that is, $\sigma(\omega)\propto|\omega|$.
IV.2.3 Localized phase